OpenStudy (anonymous):

Solve for A and B: 13x+1/2x^2+3x-2 = A/2x-1 + B/x+2 I don't know how to proceed with solving something with three unknowns.

OpenStudy (anonymous):

Here is the equation in a more readable form (see attached file)

OpenStudy (anonymous):

\[(13x + 1)/(2x ^{2} +3x - 2) = (A(x + 2) + B(2x -1) / (2x ^{2} +3x -2)\] \[13x + 1 = Ax + 2A + 2Bx - B\] 13x = Ax + 2Bx 13 = A + 2B 1 = 2A -B solve last two equation for A and B

OpenStudy (anonymous):

The answer for A = 3 and B = 5. That does not give me 3 or 5.

OpenStudy (anonymous):

It does. B = 2A -1 13 = A + 2(2A - 1) 13 = A + 4A - 2 13+ 2 = 5 A 15 = 5 A A = 3

OpenStudy (anonymous):

there is a fairly snappy way to do this

OpenStudy (anonymous):

\[\frac{13x+1}{(2x-1)(x+2)}\] to find B replace \(x\) by \(-2\) and cover up that factor in the denominator \[B=\frac{13\times -2+1}{2\times -2-1}=\frac{-25}{-5}=5\]

OpenStudy (anonymous):

@satellite73 Your way is much better but I haven't read about it so can you please explain.

OpenStudy (anonymous):

I followed you, sangya21, to this point: 13x+1 = Ax + 2A + B2x - B But I don't get how you got rid of 2A and -B from the equation.

OpenStudy (anonymous):

I just equated constant at one side and variables at other 13x = Ax + 2Bx ( equating x variables on LHS to RHS) 1 = 2A - B ( equating constants)

OpenStudy (loser66):